# Effective Bertini theorems and zeros of $p$-adic forms of degree 7

### v1

@@ 3. Facts about degree 7 forms over $\mathbb{F}_{q}$
If $L_{1},L_{2},L_{3}$ do not meet at a common point, then they contain no $\mathbb{F}_{q}$-points. Again considering the line $\langle u,v\rangle$, we have that at most two of of the $\mathbb{F}_{q^{3}}$-points $\langle u,v\rangle\cap L_{i}$ can coincide. Thus $f|_{\langle u,v\rangle}$ vanishes to order 2 on $u$ and $v$ and order 1 on at least one of the $\langle u,v\rangle\cap L_{i}$, so we may take $u_{1}=u$ and $u_{2}=v$. ∎
@@ 4.2. Consequences
Our immediate goal, accomplished below in Proposition 4.5, is show that either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or we can produce two points $u,v\in X_{f}(\mathbb{F}_{q})$ such that $f$ possesses a nonsingular geometric point on the line between them. Later, in the proof of Theorem C, we will use this in conjunction with Theorem 2.6 to produce a planar slice of $X_{f}$ with at least two (possibly singular) $\mathbb{F}_{q}$-points. The presence of these two $\mathbb{F}_{q}$-points is key to improving on [Woo08, Theorem 1].
As in [LY96], we will make use of a linear subvariety $U\subset X_{f}$ of maximal dimension. In the following lemma, we show that not only can we find points $v,w\in(X_{f}-U)(\mathbb{F}_{q})$, but we can ensure the line $\langle v,w\rangle$ is skew to $U$. This is accomplished by strengthening an argument presented in [LY96, Proof of Proposition 5.4].
@@
Let q>7q>{7} and suppose $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ is a reduced degree 7 form for $n\geq 49$. Suppose $U\subset X_{f}$ is a linear subvariety defined over $\mathbb{F}_{q}$ of maximal dimension. Then either $X_{f}(F_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $v,w\in X_{f}(\mathbb{F}_{q})$ such that
@@
We have $U(\mathbb{F}_{q})=\frac{q^{s+1}-1}{q-1}<X_{f}(\mathbb{F}_{q})$ by Proposition 4.3(iii), so there exists $v\in X_{f}(\mathbb{F}_{q})-U(\mathbb{F}_{q})$. Moreover, we have $\langle U,v\rangle\not\subset X_{f}$ by maximality of $U$, and since q>7q>{7} an application of Lemma 3.1 reveals that $f$ does not vanish on all $\mathbb{F}_{q}$-points of $\langle U,v\rangle$. Since $\langle U,v\rangle(\mathbb{F}_{q})=\frac{q^{s+2}-1}{q-1}$, Proposition 4.3(iii) again ensures there exists a point $w\in(X_{f}-\langle U,v\rangle)(\mathbb{F}_{q})$.
@@
Let q>591q>{591} and suppose $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ is a reduced degree 7 form for $n\geq 49$. Then either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $u_{1},u_{2}\in X_{f}(\mathbb{F}_{q})$ such that $\langle u_{1},u_{2}\rangle\not\subset X_{f}$ and $X_{f}\cap\langle u_{1},u_{2}\rangle$ contains a geometric point of multiplicity 1.
@@
Let $U\subset X_{f}$ be a linear subvariety defined over $\mathbb{F}_{q}$ of maximal dimension, denoted by $s$. We have $s\geq 0$ by Proposition 4.3(iii). We treat the $s=0$, $s=1$, and $s\geq 2$ cases separately.
Case $s=0$. If $s=0$, then by Proposition 4.3(iii) we have $X_{f}(\mathbb{F}_{q})\geq q+1>3$, so there exist at least four $\mathbb{F}_{q}$-points on $X_{f}$. If any four of them are colinear, then at least one has multiplicity 1, so we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. Thus there exist three points $u,v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle u,v,w\rangle\simeq\mathbb{P}^{2}$. Applying Lemma 3.7 to $f|_{\langle u,v,w\rangle}$, we have the desired result.
@@
Thus for all but at most 6 points $u\in U(\mathbb{F}_{q})$, we have that $f$ does not vanish on any lines between $u$ and a point on $(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})\supset\{v,w\}$. Instead, $f$ must vanish on a line containing either $v$, $w$, or the potential third point $x$ on $(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})$. When q>17q>{17}, this guarantees the existence of at least five points $u_{1},\ldots,u_{5}\in U(\mathbb{F}_{q})$ such that $f$ vanishes on a line $V_{i}\subset\langle u_{i},v,w\rangle$ through a common point $v$, $w$, or $x$. Without loss of generality, assume $v\in V_{i}$ for $i=1,\ldots,5$. Note that $V_{i}\cap U=\emptyset$; if not we find that $w\in\langle U,v\rangle$, contradicting our choice of $\langle v,w\rangle$ skew to $U$.
Suppose for $1\leq i<j<k\leq 5$ we have that the lines $V_{i}$, $V_{j}$, $V_{k}$ are coplanar. Recalling that $V_{i}\cap U=\emptyset$, we have that the plane $\langle V_{i},V_{j},V_{k}\rangle$ meets $U$ at a point not contained in the union of $V_{i}$, $V_{j}$, and $V_{k}$. Examining Lemma 3.5, we have that $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
Assume now that for no $(i,j,k)$ triples are $V_{i},V_{j},V_{k}$ coplanar, and set $u_{ij}=U\cap\langle V_{i},V_{j}\rangle$ for all $i<j$. The points $u_{12},u_{13},u_{23}$ are distinct, since otherwise $V_{1},V_{2},V_{3}$ are coplanar, so $f$ must not vanish on the line between one of them and $v$, by Lemma 3.6. Without loss of generality, assume $\langle u_{12},v\rangle\not\subset X_{f}$.
Consider the plane spanned by $V_{1},V_{2}$. Here $f|_{\langle V_{1},V_{2}\rangle}$ vanishes on the two lines $V_{1}$ and $V_{2}$, as well as a point $u_{12}$. By Lemma 3.5, either we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$, or $f|_{\langle V_{1},V_{2}\rangle}$ vanishes on a line defined over $\overline{\mathbb{F}_{q}}$ containing $u_{12}$. In the latter case, if the line is not defined over $\mathbb{F}_{q}$, then in fact there exist two or three conjugate lines of multiplicity 1, intersecting at $u_{12}$. This allows us to choose distinct points $u_{i}\in V_{i}(\mathbb{F}_{q})-\{v\}$ for $i=1,2$ such that $u_{12}\notin\langle u_{1},u_{2}\rangle$. Since $\langle u_{1},u_{2}\rangle$ meets the aforementioned conjugate lines in distinct $\overline{\mathbb{F}_{q}}$-points, this satisfies the desired conditions.
Finally, suppose that $u_{12}$ is contained in a line $L\subset X_{f}\cap\langle V_{1},V_{2}\rangle$ defined over $\mathbb{F}_{q}$. We have $L\cap V_{i}=\emptyset$ for $i\neq 1,2$, since no three $V_{i}$ are coplanar and $v\notin L$. Thus for $i=3,4,5$, each $V_{i}$ meets the plane $\langle L,U\rangle$ at a single point $v_{i}$; moreover, these three points are not colinear, since the $V_{i}$ were not coplanar. Thus in the plane $\langle L,U\rangle$, $f|_{\langle L,U\rangle}$ vanishes on two lines and three non-colinear points, so by Lemma 3.5 we must have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$, and we are done.
Case $s\geq 2$. As in the previous case, by Lemma 4.4, there exist $v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle v,w\rangle\not\subset X_{f}$ and is skew to $U$. First, we argue there exists a chain of linear subvarieties defined over $\mathbb{F}_{q}$,

### v2

@@ 3.1. $d=7$
If $L_{1},L_{2},L_{3}$ do not meet at a common point, then they contain no $\mathbb{F}_{q}$-points. Again considering the line $\langle u,v\rangle$, we have that at most two of of the $\mathbb{F}_{q^{3}}$-points $\langle u,v\rangle\cap L_{i}$ can coincide. Thus $f|_{\langle u,v\rangle}$ vanishes to order 2 on $u$ and $v$ and order 1 on at least one of the $\langle u,v\rangle\cap L_{i}$, so we may take $y=u$ and $z=v$. ∎
We conclude by giving what amounts to a higher dimensional version of Lemma 3.6.
Proposition 3.7.
Let q>591q>{591}. Suppose $f\in\mathbb{F}_{q}[x_{0},\ldots,x_{n}]$ is a degree 7 form defined over $\mathbb{F}_{q}$. Let $U,V,W\subset X_{f}$ be distinct linear subvarieties of dimension $s$ such that $\dim U\cap V\cap W=s-1$, $\dim\langle U,V,W\rangle=s+2$, and $f|_{\langle U,V,W\rangle}$ is nonzero with no linear factors defined over $\mathbb{F}_{q}$. Then $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}=\emptyset$ or there exist $y,z\in X_{f}(\mathbb{F}_{q})$ such that $\langle y,z\rangle\not\subset X_{f}$ and $X_{f}\cap\langle y,z\rangle$ contains a geometric point of multiplicity 1.
Proof.
The $s=0$ case is Lemma 3.6. Assume $s\geq 1$. Replacing $f$ by $f|_{\langle U,V,W\rangle}$, we may assume $n=s+2\geq 3$, $\langle U,V,W\rangle=\mathbb{P}^{n}$, and $f$ has no linear factors defined over $\mathbb{F}_{q}$.
After a change of coordinates, we may assume the intersection $U\cap V\cap W$, codimension 3 in $\mathbb{P}^{n}$, is given by
$U\cap V\cap W\colon x_{n-2}=x_{n-1}=x_{n}=0.$
For a generic plane $P\subset\mathbb{P}^{n}$, we have $P\cap U\cap V\cap W=\emptyset$. Moreover, this holds for any plane $P$ given by
$P\colon\begin{cases}x_{0}=Z,\\ x_{1}=X+\beta_{1}Z,\\ x_{2}=\gamma_{2}X+\delta_{2}Y+\beta_{2}Z,\\ \vdots\\ x_{n}=\gamma_{n}X+\delta_{n}Y+\beta_{n}Z,\end{cases}\text{ satisfying }\det\begin{pmatrix}\beta_{n-2}&\gamma_{n-2}&\delta_{n-2}\\ \beta_{n-1}&\gamma_{n-1}&\delta_{n-1}\\ \beta_{n}&\gamma_{n}&\delta_{n}\end{pmatrix}\neq 0.$
When $n=3$, we set $w_{n-2}=1$ and $z_{n-2}=0$ above.
Let $\Upsilon^{\prime}\in\mathbb{F}_{q}[B_{1},\ldots,B_{n},\Gamma_{2},\ldots,\Gamma_{n}]$ be the generic determinant above,
(3.4) $\Upsilon^{\prime}=(B_{n-1}\Gamma_{n}-B_{n}\Gamma_{n-1})(B_{n-2}\Gamma_{n}-B_{n}\Gamma_{n-2})(B_{n-2}\Gamma_{n-1}-B_{n-1}\Gamma_{n-2}).$
For any $(\boldsymbol{\beta},\boldsymbol{\gamma})$ satisfying $\Upsilon^{\prime}(\beta_{1},\ldots,\beta_{n},\gamma_{2},\ldots,\gamma_{n})\neq 0$, we define $\Xi\in\mathbb{F}_{q}[\Delta_{2},\ldots,\Delta_{n}]$ by
$\Xi=\Delta_{n-2}(\beta_{n-1}\gamma_{n}-\beta_{n}\gamma_{n-1})-\Delta_{n-1}(\beta_{n-2}\gamma_{n}-\beta_{n}\gamma_{n-2})+\Delta_{n}(\beta_{n-2}\gamma_{n-1}-\beta_{n-1}\gamma_{n-2}).$
which is necessarily nonzero.
We now apply an effective Bertini argument which is essentially a minor modification of Theorem 2.2(ii). Let $\Upsilon$ be the polynomial of degree 98 constructed in Lemma 2.10. Since q>104q>{104}, there exist $(\boldsymbol{\beta},\boldsymbol{\gamma})$ such that $\Upsilon(\boldsymbol{\beta},\boldsymbol{\gamma})\Upsilon^{\prime}(\boldsymbol{\beta},\boldsymbol{\gamma})\neq 0$.
Suppose $f$ has factorization type $(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})$ over $\overline{\mathbb{F}_{q}}$. For $1\leq i\leq t$, set
$D_{i}=\begin{cases}0&d_{i}=1,\\ 1&d_{i}>1.\end{cases}$
Consider
$\Psi=\prod_{1\leq i\leq t}\prod_{1\leq j\leq d_{i}}\Psi^{(\alpha_{j,0})}_{i,D_{i}}$
as in the proof of Theorem 2.1(ii). We have
$\deg\Psi\Xi\leq 1+\sum_{1\leq i\leq t}\frac{d_{i}D_{i}}{8}\left(-D_{i}^{3}+4d_{i}D_{i}^{2}-6D_{i}^{2}+12d_{i}D_{i}-11D_{i}+8d_{i}-6\right).$
In particular, when $q>{\deg\Psi\Xi}$, there exists $\boldsymbol{\delta}\in\mathbb{F}_{q}^{n-1}$ such that $\Psi(\boldsymbol{\delta})\Xi(\boldsymbol{\delta})\neq 0$. A straightforward computation shows $\deg\Psi\Xi\leq 127$, so our hypotheses on $q$ more than suffice for this purpose.
Since $\Xi(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\neq 0$, the plane $P$ satisfies $P\cap U\cap V\cap W=\emptyset$. In particular, the points
$u=P\cap U,v=P\cap V,w=P\cap W$
are distinct, noncolinear points in $(X_{f}\cap P)(\mathbb{F}_{q})$. Moreover, since $f|_{P}$ has the same factorization type as $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}=0$, and the latter only has linear factors inherited from $f$ itself which remain distinct in $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$, any linear factors of $f|_{P}$ are defined over $\overline{\mathbb{F}_{q}}$ but not $\mathbb{F}_{q}$. Thus $f|_{P}$ satisfies the hypotheses of Lemma 3.6 and we are done. ∎
3.2. $d=11$
The same arguments in degree 11 give analogues of Lemmas 3.4, 3.5, which we use to give analogues of Lemma 3.6 and Proposition 3.7.
Lemma 3.8.
Let $q\geq{11}$. Suppose $f(x_{0},\ldots,x_{m})$ is a degree 11 form and there exist $k$ distinct linear subvarieties $U_{1},\ldots,U_{k}\subset X_{f}\subset\mathbb{P}^{n}$ of dimension $s$ such that $\dim\langle U_{1},\ldots,U_{k}\rangle=s+1$.
  1. (i)
    If $k\geq 6$ then $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or $\langle U_{1},\ldots,U_{k}\rangle\subset X_{f}$.
  2. (ii)
    If $k=5$ and there exists $v\in X_{f}\cap\langle U_{1},\ldots,U_{5}\rangle$ such that $v\notin U_{1},\ldots,U_{5}$, then $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or $\langle U_{1},\ldots,U_{5}\rangle\subset X_{f}$.
Lemma 3.9.
Let $q\geq{11}$. Suppose $U\subset X_{f}$ is a linear subvariety of positive dimension defined over $\mathbb{F}_{q}$, and $v\in X_{f}(\mathbb{F}_{q})-U$. If there exist distinct codimension 1 linear subvarieties $U_{1}^{(1)},\ldots,U_{5}^{(1)}\subset U$ such that
$\langle U_{i}^{(1)},v\rangle\subset X_{f}\text{ for }i=1,\ldots,5,$
Lemma 3.10.
Let q>7061q>{7061}. Suppose $f\in\mathbb{F}_{q}[x_{0},x_{1},x_{2}]$ is a nonzero ternary degree 11 form with no linear factors defined over $\mathbb{F}_{q}$. Suppose further that there exist $u,v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle u,v,w\rangle=\mathbb{P}^{2}$. Then either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $y,z\in X_{f}(\mathbb{F}_{q})$ such that $\langle y,z\rangle\not\subset X_{f}$ and $X_{f}\cap\langle y,z\rangle$ contains a geometric point of multiplicity 1.
Proof.
The possible factorization types of $f$ over $\mathbb{F}_{q}$ are
$(11),(9,2),(8,3),(7,4),(7,2,2),(7,2^{2}),(6,5),(6,3,2),$
$(5,4,2),(5,3,3),(5,3^{2}),(5,2,2,2),(5,2^{2},2),(5,2^{3}),(4,4,3),(4^{2},3),(4,3,2,2),(4,3,2^{2}),$
$(3,3,3,2),(3^{2},3,2),(3^{3},2),(3,2,2,2,2),(3,2^{2},2,2),(3,2^{2},2^{2}),(3,2^{3},2),\text{ or }(3,2^{4}).$
If any of the factors of multiplicity 1 defined over $\mathbb{F}_{q}$ are irreducible over $\overline{\mathbb{F}_{q}}$, then we can use Lemma 3.2 to find $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$, as in the proof of Lemma 3.6. The limiting case is that of $f$ geometrically irreducible, where the presence of the three singular $\mathbb{F}_{q}$-points $u,v,w$ means q>7061q>{7061} suffices.
Write the factorization over $\mathbb{F}_{q}$ as $f=\prod_{1\leq i\leq t}g_{i}^{e_{i}}$, i.e. if $f$ has type $(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})$ over $\mathbb{F}_{q}$, then $g_{i}$ is defined over $\mathbb{F}_{q}$ and has degree $d_{i}$. Assume for all $i$ with $e_{i}=1$, we have $g_{i}$ is geometrically reducible; otherwise we are in the situation above.
Suppose there exists an $i$ with the following properties:
  1. (i)
    $d_{i}$ is prime,
  2. (ii)
    $e_{i}=1$,
  3. (iii)
    $d_{i}>\frac{d_{j}}{2}$ for all $j\neq i$ with $e_{j}=1$,
  4. (iv)
    $d_{i}>d_{j}$ for all $j\neq i$ with $e_{j}>1$.
Since $g_{i}$ is reducible over $\overline{\mathbb{F}_{q}}$, it must factor into $d_{i}$ linear forms defined over $\mathbb{F}_{q^{d_{i}}}$, cyclically permuted by the action of $\mathrm{Gal}(\mathbb{F}_{q^{d_{i}}}/\mathbb{F}_{q})$. In particular, $g_{i}$ vanishes on at most one $\mathbb{F}_{q}$-point, so without loss of generality it vanishes on $d_{i}$ distinct points on the line $\langle u,v\rangle$ defined (and conjugate) over $\mathbb{F}_{q^{d_{i}}}$.
Consider now how the $g_{j}$ vanish on $\langle u,v\rangle$ for $j\neq i$. If $e_{j}=1$, then $g_{j}$ — and its restriction to the line $\langle u,v\rangle$ — factors into forms of degree at most $d_{j}/2$ over $\overline{\mathbb{F}_{q}}$. By (iii), the degree of these factors of $g_{j}$ is too small to vanish on any $\mathbb{F}_{q^{d_{i}}}$-points of the line. The same argument works for $e_{j}>1$, since we have the stronger hypothesis (iv). Thus $u,v\in X_{f}(\mathbb{F}_{q})$ satisfy the desired requirements, with the zeros of $g_{i}$ on this line providing the necessary geometric point(s) of multiplicity 1.
Conditions (i) – (iv) are satisfied for all types except $(11)$, $(9,2)$, $(8,3)$, $(6,3,2)$, $(4^{2},3)$, $(3^{2},3,2)$, and $(3^{3},2)$. We deal with the stragglers ad hoc.
  • $(11)$
    If $f$ is geometrically reducible, then it must factor into 11 conjugate linear forms defined over $\mathbb{F}_{q^{11}}$. However, this is not possible, as such $f$ vanishes on at most one $\mathbb{F}_{q}$-point.
  • $(9,2)$
    If $g_{1}$ is not geometrically irreducible, it must factor into three conjugate cubic forms $h,h^{\prime},h^{\prime\prime}$ over $\mathbb{F}_{q^{3}}$. As above, we may assume the quadratic factor $g_{2}$ vanishes on two distinct $\mathbb{F}_{q^{2}}$-points of $\langle u,v\rangle$. Since $g_{1}$ must then vanish on both $u,v\in X_{f}(\mathbb{F}_{q})$, each of $h,h^{\prime},h^{\prime\prime}$ must vanish on both $u$ and $v$, and hence at most one more $\mathbb{F}_{q}$-point on $\langle u,v\rangle$, so the line $\langle u,v\rangle$ satisfies the desired requirements.
  • $(8,3)$
    If $g_{1}$ is not geometrically irreducible, it must factor into two conjugate quartic forms $h,h^{\prime}$ over $\mathbb{F}_{q^{2}}$, while we may assume the cubic factor $g_{2}$ vanishes on three distinct conjugate $\mathbb{F}_{q^{3}}$-points of $\langle u,v\rangle$. As in the $(9,2)$ case, $h,h^{\prime}$ are forced to vanish on $u,v$, so they cannot vanish on any $\mathbb{F}_{q^{3}}$-points along $\langle u,v\rangle$.
  • $(6,3,2)$
    Here $g_{1}$ can factor in two ways. Suppose it factors into conjugate cubic forms $h,h^{\prime}$ defined over $\mathbb{F}_{q^{2}}$ and assume the cubic factor $g_{2}$ vanishes on three distinct $\mathbb{F}_{q^{3}}$-points of $\langle u,v\rangle$. The quadratic factor $g_{3}$ vanishes on at most one of $u$ or $v$, thus $h,h^{\prime}$ must both vanish on (at least) the other. Their degrees are then too small to also vanish on any $\mathbb{F}_{q^{6}}$-points on $\langle u,v\rangle$, nor in particular any $\mathbb{F}_{q^{3}}$-points, so we are done.
    If instead $g_{1}$ factors into conjugate quadratic forms $h,h^{\prime},h^{\prime\prime}$ defined over $\mathbb{F}_{q^{3}}$, a similar argument shows they all must vanish on at least one of $u$ or $v$ and one other $\mathbb{F}_{q}$-point on the line $\langle u,v\rangle$, while the quadratic factor $g_{3}$ vanishes on distinct $\mathbb{F}_{q^{2}}$-points of $\langle u,v\rangle$. The cubic factor $g_{2}$ meets the line at an $\mathbb{F}_{q}$-point or distinct $\mathbb{F}_{q^{3}}$-points, but eitherway the requirements of the lemma are met.
  • $(4^{2},3)$
    Assume the cubic factor $g_{2}$ meets the line $\langle u,v\rangle$ at three distinct $\mathbb{F}_{q^{3}}$-points. For degree reasons only, the quartic factor $g_{1}$ can vanish on at most two of them, since it must vanish on $u$ and $v$.
  • $(3^{2},3,2)$
    Suppose the cubic factor $g_{1}$ is irreducible over $\overline{\mathbb{F}_{q}}$. The cubic factor $g_{2}$ vanishes on at most one $\mathbb{F}_{q}$-point in the plane. Let $y,z\in X_{g_{1}}(\mathbb{F}_{q})$ be distinct points such that $\langle y,z\rangle$ avoids this point (if it exists); note that the hypotheses on $q$ ensure more than enough $\mathbb{F}_{q}$-solutions to $g_{1}=0$ for this. On the line $\langle y,z\rangle$, $g_{1}$ vanishes on $y$, $z$, and at least one more point, necessarily defined over $\mathbb{F}_{q}$; the cubic factor $g_{2}$ vanishes in three distinct $\mathbb{F}_{q^{3}}$-points, while the quadratic $g_{3}$ vanishes on an $\mathbb{F}_{q}$-point or two $\mathbb{F}_{q^{2}}$-points. In either case, the zeros of $g_{2}$ along this line provide the necessary points of multiplicity 1.
    If $g_{1}$ is reducible over $\overline{\mathbb{F}_{q}}$, then in fact $X_{f}(\mathbb{F}_{q})=\{u,v,w\}$, with each of $g_{1},g_{2},g_{3}$ vanishing on exactly one. If $u\in X_{g_{1}}$, then either of $\langle u,v\rangle$ or $\langle u,w\rangle$ satisfy the desired requirements.
  • $(3^{3},2)$
    Assume the quadratic factor $g_{2}$ meets the line $\langle u,v\rangle$ at two distinct $\mathbb{F}_{q^{2}}$-points. As in the $(4^{2},3)$ case, the cubic factor $g_{1}$ can vanish on at most one of them, since it must vanish on $u$ and $v$.
Proposition 3.11.
Let q>7061q>{7061}. Suppose $f\in\mathbb{F}_{q}[x_{0},\ldots,x_{n}]$ is a degree 11 form defined over $\mathbb{F}_{q}$. Let $U,V,W\subset X_{f}$ be distinct linear subvarieties of dimension $s$ such that $\dim U\cap V\cap W=s-1$, $\dim\langle U,V,W\rangle=s+2$, and $f|_{\langle U,V,W\rangle}$ is nonzero with no linear factors defined over $\mathbb{F}_{q}$. Then $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}=\emptyset$ or there exist $y,z\in X_{f}(\mathbb{F}_{q})$ such that $\langle y,z\rangle\not\subset X_{f}$ and $X_{f}\cap\langle y,z\rangle$ contains a geometric point of multiplicity 1.
Proof.
The proof is essentially identical to that of Proposition 3.7, except that we reduce to Lemma 3.10 instead of Lemma 3.6. This accounts for the hypothesis q>7061q>{7061}. We briefly sketch the other differences.
The polynomial $\Upsilon$ from Lemma 2.10 now has degree at most 242, while $\Upsilon^{\prime}$ is the same as in the proof of Proposition 3.7. Thus there exists $(\boldsymbol{\beta},\boldsymbol{\gamma})\in\mathbb{F}_{q}^{2m-2}$ for which $\Upsilon\Upsilon^{\prime}\neq 0$.
With $\Psi$ defined the same way as in the proof of Proposition 3.7, its degree is at most 330 by Theorem 2.9. The polynomial $\Xi$ is again linear. Thus $q$ is sufficiently large to guarantee the existence of $\boldsymbol{\delta}$ with $\Psi\Xi\neq 0$.
Finally, $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has no linear factors defined over $\mathbb{F}_{q}$, corresponding to a plane $P$ not containing $U\cap V\cap W$ such that $X_{f}\cap P$ has no linear components defined over $\mathbb{F}_{q}$. Since $P$ meets $U,V,W$ in distinct noncolinear points $u,v,w$, the hypotheses of Lemma 3.10 are satisfied, completing the proof. ∎
@@ 4.2. Consequences
Our immediate goal, accomplished below in Proposition 4.5, is show that either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or we can produce two points $x,y\in X_{f}(\mathbb{F}_{q})$ such that $f$ possesses a nonsingular geometric point on the line between them. Later, in the proof of Theorem B, we will use this in conjunction with Theorem 2.9 to produce a planar slice of $X_{f}$ with at least two (possibly singular) $\mathbb{F}_{q}$-points. The presence of these two $\mathbb{F}_{q}$-points is key to improving on [WOO08, Theorem 1].
As in [LY96], we will make use of a linear subvariety $U\subset X_{f}$ of maximal dimension. In the following lemma, we show that not only can we find points $v,w\in(X_{f}-U)(\mathbb{F}_{q})$, but we can ensure the line $\langle v,w\rangle$ is skew to $U$. This is accomplished by strengthening an argument presented in [LY96, Proof of Prop. 5.4].
@@
Let q>dq>{d} and suppose $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ is a reduced degree $d$ form for $n\geq d^{2}$. Suppose $U\subset X_{f}$ is a linear subvariety defined over $\mathbb{F}_{q}$ of maximal dimension. Then either $X_{f}(F_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $v,w\in X_{f}(\mathbb{F}_{q})$ such that
@@
Let $s=\dim U$. We have $U(\mathbb{F}_{q})=\frac{q^{s+1}-1}{q-1}<X_{f}(\mathbb{F}_{q})$ by Proposition 4.3(iii), so there exists $v\in X_{f}(\mathbb{F}_{q})-U(\mathbb{F}_{q})$. Moreover, we have $\langle U,v\rangle\not\subset X_{f}$ by maximality of $U$, and since q>dq>{d} an application of Lemma 3.1 reveals that $f$ does not vanish on all $\mathbb{F}_{q}$-points of $\langle U,v\rangle$. Since $\langle U,v\rangle(\mathbb{F}_{q})=\frac{q^{s+2}-1}{q-1}$, Proposition 4.3(iii) again ensures there exists a point $w\in(X_{f}-\langle U,v\rangle)(\mathbb{F}_{q})$.
@@
Let $d=7$ and $n\geq 49$. Suppose q>591q>{591} and $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ is a reduced form of degree $7$. Then either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $y,z\in X_{f}(\mathbb{F}_{q})$ such that $\langle y,z\rangle\not\subset X_{f}$ and $X_{f}\cap\langle y,z\rangle$ contains a geometric point of multiplicity 1.
@@
Our approach is to find noncolinear points $u,v,w\in X_{f}(\mathbb{F}_{q})$ such that $f|_{\langle u,v,w\rangle}$ is nonzero and has no linear factors over $\mathbb{F}_{q}$. Then we can apply Lemma 3.6. Let $U\subset X_{f}$ be a linear subvariety defined over $\mathbb{F}_{q}$ of maximal dimension, denoted by $s$. We have $s\geq 0$ by Proposition 4.3(iii). We treat the $s=0$, $s=1$, and $s\geq 2$ cases separately.
Case $s=0$. If $s=0$, then by Proposition 4.3(iii) we have $X_{f}(\mathbb{F}_{q})\geq q+1$, so there exist at least four $\mathbb{F}_{q}$-points on $X_{f}$. If any four lie on a line, then $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ by Lemma 3.4. Otherwise, there exist three points $u,v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle u,v,w\rangle\simeq\mathbb{P}^{2}$. Applying Lemma 3.6 to $f|_{\langle u,v,w\rangle}$, we have the result.
@@
Thus for all but at most 6 points $u\in U(\mathbb{F}_{q})$, we have that $f$ does not vanish on any lines between $u$ and a point on $(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})\supset\{v,w\}$. Instead, $f$ must vanish on a line containing either $v$, $w$, or the potential third point $x$ on $(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})$. When q>11q>{11}, this guarantees the existence of at least three points $u_{1},\ldots,u_{3}\in U(\mathbb{F}_{q})$ such that $f$ vanishes on a line $V_{i}\subset\langle u_{i},v,w\rangle$ through a common point $v$, $w$, or $x$. Without loss of generality, assume $v\in V_{i}$ for $i=1,\ldots,3$. Note that $V_{i}\cap U=\emptyset$; if not we find that $w\in\langle U,v\rangle$, contradicting our choice of $\langle v,w\rangle$ skew to $U$.
Suppose the lines $V_{1},V_{2},V_{3}$ are coplanar. Recalling that $V_{i}\cap U=\emptyset$, we have that the plane $\langle V_{1},V_{2},V_{3}\rangle$ meets $U$ at a point not contained in the union of $V_{1}$, $V_{2}$, and $V_{3}$. By Lemma 3.4(ii), $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
Suppose now that $V_{1},V_{2},V_{3}$ are not coplanar, hence $\dim\langle V_{1},V_{2},V_{3}\rangle=3$. Thus they satisfy the hypotheses of Proposition 3.7, and we reach the desired conclusion.
Case $s\geq 2$. By Lemma 4.4, there exist $v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle v,w\rangle\not\subset X_{f}$ and is skew to $U$. First, we argue there exists a chain of linear subvarieties defined over $\mathbb{F}_{q}$,